OpenStudy (tal_tal):

Fan and medal! The ratio of the number of cups of yellow paint to the number of cups of blue paint that Greg mixed to make a shade of green paint is 2:1. Part A: Write a sentence to describe the number of cups of yellow paint that Greg mixed with each cup of blue paint. (4 points) Part B: How many cups of blue paint should Greg add to each cup of yellow paint to make the shade of green? Explain your reasoning. (4 points) Part C: If Greg purchased 6 cups of yellow paint for $54, what is the unit price per cup of yellow paint? (2 points)

OpenStudy (tal_tal):

@bibby

OpenStudy (tal_tal):

I know it's a lot but I'm really stuck

OpenStudy (bibby):

the unit price is a simplified fraction. if we want the unit price of price per cup we set up a fraction such that we have the price over the amount of cups \(\dfrac{$54}{6 \text{ cups}}=\dfrac{$9}{1cup}\)

OpenStudy (tal_tal):

Which part is that for

OpenStudy (bibby):

I said unit price there's only one part that asked for unit price

OpenStudy (tal_tal):

oh okay

OpenStudy (bibby):

2:1 yellow:blue part a: There are twice the amount of yellow cups of paint as there are blue cups of paint.

OpenStudy (bibby):

part b: since we want 2 yellow + 1 blue 2y+1b we wanna know how much blue gets added to 1 cup of yellow so we divide both sides by 2 y+\(\dfrac{1}{2}b\)

OpenStudy (tal_tal):

Thank you!